# Theoretical Detailed Analyses for DC readout and a Fabri-Pérot gravitational-wave detector

### v1

@@ VIII.2 Incomplete equilibrium tuning
The behavior of the noise-spectral density $S_{H}(\Omega,\theta=\pi/4)$ in Eq. (298) is depicted in Fig. 11. Since the effects of the incomplete equilibrium tuning are factorized $1+\epsilon^{2}{{\mathfrak{R}}}^{2}]$, the shape of the noise spectral densities does not show any drastic changes. However, we can see that if the effect of the additional classical carrier $\epsilon{{\mathfrak{R}}}$ is considered, the power dependence of the noise spectral density is changed. As a result, the bottom figures in Fig. 11 show the property that the shot noise is not reduced even if the input power $I_{0}$ is increased. Incidentally, the bottom right panel in Fig. 11 is the same figure as Fig. 6.
Finally, we discuss the realization of the ideal noise spectral density $S_{H}(\Omega)$ with $\epsilon{{\mathfrak{R}}}=0$. As shown in both Fig. 10 and Fig. 11, if we have sufficiently small $\epsilon$, we can realize the ideal noise spectral density $S_{H}(\Omega)$ with $\epsilon{{\mathfrak{R}}}\sim 0$. From the input-output relation (189), if the term $\mathfrak{R}$}} in the classical carrier field, which is proportional to $2\pi\delta(\Omega)$ is negligible compared with $1$ in the same term, we can realize the ideal input-output relation. In the case of the incomplete equilibrium tuning, the phenomenological parameter $\epsilon$, which we introduced, appears as $\epsilon{{\mathfrak{R}}}$ in this classical carrier field of the input-output relation as mentioned above. Therefore, if we achieve the $\epsilon$ so that
$\epsilon{{\mathfrak{R}}}\lesssim O(10^{-1}),$ (299)
we can realize the idealized input-output relation which as shown in Ref. [8]. From the order of magnitude (218) of $\mathfrak{R}$}}, $\epsilon$ can be estimated as
$\epsilon$ $\lesssim$ $O(10^{-1}){{\mathfrak{R}}}^{-1}$ (301)
$\sim$ $O(10^{-1})\frac{mc^{2}\omega_{p}^{2}T^{2}}{64I_{0}\omega_{0}}$
$\sim$ $3\times 10^{-4}\times\left(\frac{10^{2}W}{I_{0}}\right)\left(\frac{2\pi\times 10^{14}\mbox{Hz}}{\omega_{0}}\right)\left(\frac{m}{40\mbox{kg}}\right)$
@@
$\epsilon{{\mathscr{D}}}_{EM}$ $\lesssim$ $+O(10^{-1})\frac{cT}{16\omega_{0}}$
$\sim$ $+3\times 10^{-11}\;\mbox{m}\times\left(\frac{T}{10^{-2}}\right)\left(\frac{2\pi\times 10^{14}\;\mbox{Hz}}{\omega_{0}}\right)$
@@
$\epsilon{{\mathscr{D}}}_{ITM}$ $\lesssim$ $-O(10^{-1})\frac{cT}{16\omega_{0}}$
$\sim$ $-3\times 10^{-11}\;\mbox{m}\times\left(\frac{T}{10^{-2}}\right)\left(\frac{2\pi\times 10^{14}\;\mbox{Hz}}{\omega_{0}}\right).$
@@ IX Summary and Discussion
The primary purpose of this paper was to develop the theoretical description of the quantum noise in gravitational-wave detectors in detail, which was motivated by the theoretical developments of the measurement theory in quantum field theories. In the mathematically rigorous measurement theory in quantum mechanics [12], the specification of the finally observed quantum operator is an essential issue in the theory. We must extend this mathematical theory to quantum field theories to apply it to gravitational-wave detectors. Besides this extension of the quantum measurement theory to quantum field theories, the above specification of the finally observed quantum operator is crucially important. Actually, in quantum measurement theories, there is a famous problem of the von Neumann chain [17], i.e., where we should regard the measurement outcomes as the classical information in the sequence of the quantum measurements as emphasized in Sec. I. In this paper, we adopt the standing point that the von Neumann chain should be cut at the moment that the photodetectors convert the information in photon signals to the electric currents. Note that actual gravitational-wave detectors employ a feedback system to control the stability of the detector. However, from our standpoint, we regard the feedback current as a classical electric current. Therefore, we do not have to care about the quantum properties of the feedback control system. If this premise is wrong, we have to discuss quantum feedback control systems [18] as emphasized in Sec. I, which is beyond the current scope of this paper.
@@
After the introduction of the notions which was used within this paper in Sec. II, we first developed the description of the DC readout scheme in terms of quantum electrodynamics. Although the analysis in these arguments of the DC readout scheme is just an extension of the balanced homodyne detection, which was discussed in Ref. [19], the essence of the DC readout scheme is the leakage of the large classical carrier field from the output port. Due to this large classical carrier field, we can separate the output electric field into orders of the large classical carrier field. After excluding the first dominant term, we can measure the output field, which includes gravitational signals as the second dominant term. Since the leaked large classical carrier field from the output port serves as the reference for the output signal in the DC readout scheme, the leaked large classical carrier field is essential in the DC readout scheme.
@@
After the general arguments of the DC readout scheme in Sec. III, we carefully discussed the input-output relation of the Fabri-Pérot gravitational-wave detectors, because the details of the input-output relation are necessary to examine the properties of the DC readout scheme. Therefore, in Secs. IV, V, and VI, we described the detailed theoretical description of the derivation of the input-output relation of the Fabri-Pérot interferometer.
@@
In Sec. IV, we derived the input-output relation through the propagation effects and the reflection effects due to the mirrors in the Fabri-Pérot interferometer. In the previous works of gravitational-wave detectors, it has been considered only the end mirrors’ motions. However, in this paper, we also take into account the motions of the intermediate mirrors. We obtained the input-output relation without any specification of the mirror displacements. The equations of motion of mirrors determine these mirror displacements. Since we discussed quantum motions of mirrors, we used the Heisenberg equations of motion in quantum mechanics as the equation of motion for mirrors. Furthermore, we introduce the fundamental frequency $\omega_{p}$ of the mirrors’ pendulum, where $\omega_{p}$ is estimated as $\sim 1$Hz.
@@
In Sec. V, we briefly reviewed a quantum mechanical forced harmonic oscillator, which is essential to our arguments on the motions of mirrors in gravitational-wave detectors. Based on the understanding of this quantum forced harmonic oscillator, we explicitly derived the equations of motion for the mirrors in the Fabri-Përot interferometer in Sec. V.2. Then, we pointed out that the evaluation of the radiation pressure effects is essential. Although the details are described in Appendix B, we review the outline of the derivation of the radiation pressure forces to the intermediate and end mirrors in Fabri-Pérot gravitational-wave detectors in Sec. V.3. From the derived radiation pressure forces, we obtained the solutions to the Heisenberg equations in Sec. V.4.
@@
Together with the input-output relation derived in Sec. IV and the solutions to the Heisenberg equation derived in Sec. V.4, we obtain the final input-output relation of the Fabri-Pérot interferometer in Sec. VI. Then, we reach the input-output relation for the Fabri-Pérot interferometer (189). Although the parameters $\beta$, $\kappa$, and $h_{SQL}$ in the input-output relation are different from these parameters in Ref. [8], the resulting input-output relation is almost the same as that derived in Ref. [8]. The difference between $\kappa$ in this paper and the $\kappa$ in Ref. [8], which denote $\kappa^{(K)}$ in this paper are depicted in Fig. 3. The difference between $h_{SQL}^{(EQ)}$ ($m=m_{EM}=m_{ITM}$) and $h_{SQL}$ in Ref. [8] which denote $h_{SQL}^{(K)}$ within this paper are depicted in Fig. 4. The frequency dependence of these differences comes from the phase difference between the motion of EMs and the motion of ITMs. Due to these differences, we reach the difference of the $S_{(K)}^{(EQ)}$ and $S_{(K)}^{(K)}$ as depicted in Fig. 5, where $S_{(K)}^{(EQ)}$ is the equal mass version of Eq. (197) and $S_{(K)}^{(K)}$ is defined by Eq. (212).
@@
In addition to the above difference between the input-output relation (189), the difference also exists in the classical carrier part as the parameter $\mathfrak{R}$}}. This classical carrier field with the modification $\mathfrak{R}$}} is used as the reference field to measure the gravitational-wave signal in the DC readout scheme. Therefore, the classical carrier field is essential in the DC readout scheme. As mentioned in Sec. VII, the expectation value of the signal is not affected by the additional parameter $\mathfrak{R}$}} in the classical carrier field due to the reality condition of the gravitational-wave signal $H(\Omega)=H^{*}(-\Omega)$. The contribution of $\mathfrak{R}$}} in the classical carrier field affects the signal-referred noise spectral density (234). For this reason, we discussed the difference from the Kimble noise spectral density in Ref. [8]. As a result of the additional classical carrier field $\mathfrak{R}$}}, we have to conclude that the shot noise in the high-frequency range does not decrease even if the incident laser power is increased. This contradicts to the description in Ref. [8] which is the common knowledge in the community of gravitational-wave experiments.
@@
In this situation, we reconsidered the tuning points (161) and (177) in Sec. VIII. The origin of the additional modification $\mathfrak{R}$}} of the classical carrier in the input-output relation (189) is the leakage of the classical carrier from the radiation pressure forces, which are proportional to $2\pi\delta(\Omega)$ in the Fourier transformation of the radiation pressure forces. In the time domain, these terms, proportional to $2\pi\delta(\Omega)$, represent the constant forces in the radiation pressure forces. Here, we note that the equations of motion for mirrors are given in the form of a forced harmonic oscillator. In simple classical mechanics, or even in quantum mechanics, a constant force in a harmonic oscillator yields the deviation of the equilibrium point. If there exists a constant force in the Heisenberg equation of motion for a harmonic oscillator, we have to take into account the change of the equilibrium points of the mirror displacements. For this reason, we reconsidered the Heisenberg equation of motion for the mirrors and the tuning points (161) and (177) in Sec. VIII.
@@
We also considered the incomplete equilibrium tuning in which the elimination of the deviations of the mirror displacements from their equilibrium points is not complete through the introduction of the phenomenological parameter $\epsilon$. We call this elimination the “incomplete equilibrium tuning.” If $\epsilon=0$, we can realize the complete equilibrium tuning discussed above. Instead of the complete equilibrium tuning (261) and (262), we apply the tuning condition (288) and (289). We also showed that the additional modification $\mathfrak{R}$}} of the classical carrier in the input-output relation is explicitly expressed by the displacement of the equilibrium points ${{\mathscr{D}}}_{EM}$ and ${{\mathscr{D}}}_{ITM}$. In this incomplete equilibrium tuning, the stationary noise spectral density is given by (292). To clarify the properties of the noise spectral density (292), we consider the typical two cases. The first case is $\theta\ll 1$ but $\theta\neq 0$ and the second case is $\theta=\pi/4$. These are depicted in Figs. 10 and 11.
In the first case, the noise spectral density is given by Eq. (294). If $\epsilon$ is sufficiently large, Fig. 10 indicates the dip in the low-frequency region. Although this dip violates the “standard quantum limit” which is the envelope of the curves with different powers $I_{0}$ in $\epsilon=0$ case, this is not surprising fact that there is no contradiction with the commutation relation $\hat{X}(t),\hat{P}(t)]=i\hbar$ as discussed in Sec. V. In the second case, the effect of $\mathfrak{R}$}} appears as the factor of the noise spectral density as Eq. (298). In both cases, if the incompleteness $\epsilon$ is sufficiently large, there are cases where the shot noise in the high-frequency region does not decrease even if the incident power is increased. We also evaluate the incompleteness $\epsilon$ when we can realize the behavior near the complete equilibrium tuning in Eqs. (LABEL:eq:classical-terms-of-RPF-EM-renorm-retune-require-leading-order) and (LABEL:eq:classical-terms-of-RPF-ITM-renorm-retune-require-leading-order). This is the main result of this paper.
Here, we note that we cannot discuss the recent experimental results of the violation of the “standard quantum limit” [10, 11] through the ingredients of this paper, because we did not discuss the power recycling, the signal recycling, or the squeezed input techniques. We cannot say any relation between the dips depicted in Fig. 10 and works in Refs. [10, 11] due to the same reason. In other words, there are many rooms to be developed through an accurate understanding within this paper. We leave these rooms for future work.
@@
Finally, we have to emphasize that the purpose of this paper is not to point out new techniques nor to develop new ideas for gravitational-wave detectors themselves. However, in this paper, we aim to explicitly show the derivation of the common knowledge in the community of the gravitational-wave detections to clarify their concepts from a more theoretically accurate point of view. Due to the more precise theoretical arguments on the noise spectral density, we could discuss the incomplete equilibrium tuning of the Frabri-Pérot interferometer as an imperfection of the interferometric gravitational-wave detectors. We hope the ingredients of this paper and this kind of research will be helpful for the further development of detector science in gravitational-wave detectors.

### v2

@@ VIII.2 Incomplete equilibrium tuning
The behavior of the noise-spectral density $S_{H}(\Omega,\theta=\pi/4)$ in Eq. (298) is depicted in Fig. 11. Since the effects of the incomplete equilibrium tuning are factorized $1+\epsilon^{2}{{\mathfrak{R}}}^{2}]$, the shape of the noise spectral densities does not show any drastic changes. However, we can see that if the effect of the additional classical carrier $\epsilon{{\mathfrak{R}}}$ is considered, the power dependence of the noise spectral density is changed. As a result, the bottom figures in Fig. 11 show the property that the shot-noise contribution is not reduced even if the input power $I_{0}$ is increased. Incidentally, the bottom right panel in Fig. 11 is the same figure as Fig. 6.
Finally, we discuss the realization of the ideal noise spectral density $S_{H}(\Omega)$ with $\epsilon{{\mathfrak{R}}}=0$. As shown in both Fig. 10 and Fig. 11, if we have sufficiently small $\epsilon$, we can realize the ideal noise spectral density $S_{H}(\Omega)$ with $\epsilon{{\mathfrak{R}}}\sim 0$. From the input-output relation (189), if the term $\mathfrak{R}$}} in the classical carrier field, which is proportional to $2\pi\delta(\Omega)$ is negligible compared with $1$ in the same term, we can realize the ideal input-output relation. In the case of the incomplete equilibrium tuning, the incomplete parameter $\epsilon$, which we introduced, appears as $\epsilon{{\mathfrak{R}}}$ in this classical carrier field of the input-output relation as mentioned above. Therefore, if we achieve the $\epsilon$ so that
$\epsilon{{\mathfrak{R}}}\lesssim O(10^{-2}),$ (299)
we can realize the idealized input-output relation which is shown in Ref. [10]. From the order of magnitude (218) of $\mathfrak{R}$}}, $\epsilon$ can be estimated as
$\epsilon$ $\lesssim$ $O(10^{-2}){{\mathfrak{R}}}^{-1}$ (301)
$\sim$ $O(10^{-2})\frac{mc^{2}\omega_{p}^{2}T^{2}}{64I_{0}\omega_{0}}$
$\sim$ $3\times 10^{-5}\times\left(\frac{10^{2}W}{I_{0}}\right)\left(\frac{2\pi\times 10^{14}\mbox{Hz}}{\omega_{0}}\right)\left(\frac{m}{40\mbox{kg}}\right)$
@@
$\epsilon{{\mathscr{D}}}_{EM}$ $\lesssim$ $+O(10^{-2})\frac{cT}{16\omega_{0}}$
$\sim$ $+3\times 10^{-12}\;\mbox{m}\times\left(\frac{T}{10^{-2}}\right)\left(\frac{2\pi\times 10^{14}\;\mbox{Hz}}{\omega_{0}}\right)$
@@
$\left|\epsilon{{\mathscr{D}}}_{ITM}\right|$ $\lesssim$ $O(10^{-2})\frac{cT}{16\omega_{0}}$
$\sim$ $3\times 10^{-12}\;\mbox{m}\times\left(\frac{T}{10^{-2}}\right)\left(\frac{2\pi\times 10^{14}\;\mbox{Hz}}{\omega_{0}}\right).$
@@ IX Summary and Discussion
The primary purpose of this paper was to develop a detailed theoretical description of quantum noise in gravitational-wave detectors, motivated by advances in measurement theory in quantum field theory. In the mathematically rigorous measurement theory in quantum mechanics [14], the specification of the final observed quantum operator is an essential issue in the theory. We must extend this mathematical theory to quantum field theories to apply it to gravitational-wave detectors. Besides this extension of quantum measurement theory to quantum field theories, the above specification of the final observed quantum operator is crucial. Actually, in quantum measurement theories, there is a famous problem of the von Neumann chain [19], i.e., where we should regard the measurement outcomes as the classical information in the sequence of the quantum measurements as emphasized in Sec. I. In this paper, we adopt the view that the von Neumann chain should be cut at the point where photodetectors convert photon signals into electric currents. Note that actual gravitational-wave detectors employ a feedback system to control the stability of the detector. However, from our standpoint, we regard the feedback current as a classical electric current. Therefore, we do not need to consider the quantum properties of the feedback control system. If this premise is wrong, we have to discuss quantum feedback control systems [20] as emphasized in Sec. I, which is beyond the current scope of this paper.
@@
After the introduction of the notions that were used within this paper in Sec. II, we first developed the description of the DC readout scheme in terms of quantum electrodynamics. Although the analysis in these arguments of the DC readout scheme is just an extension of the balanced homodyne detection, which was discussed in Ref. [21], the essence of the DC readout scheme is the leakage of the large classical carrier field from the output port. Due to this large classical carrier field, we can separate the output electric field into orders of the large classical carrier field. After excluding the first dominant term, we can measure the output field, which includes gravitational signals as the second dominant term. Since the leaked large classical carrier field from the output port serves as the reference for the output signal in the DC readout scheme, the leaked large classical carrier field is essential in the DC readout scheme.
@@
After the general arguments of the DC readout scheme in Sec. III, we carefully discussed the input-output relation of the Fabry-Pérot gravitational-wave detectors, because the details of the input-output relation are necessary to examine the properties of the DC readout scheme. Therefore, in Secs. IV, V, and VI, we described the detailed theoretical description of the derivation of the input-output relation of the Fabry-Pérot interferometer.
@@
In Sec. IV, we derived the input-output relation through the propagation effects and the reflection effects due to the mirrors in the Fabry-Pérot interferometer. In the previous works of gravitational-wave detectors, it has been considered only the end mirrors’ motions. However, in this paper, we also account for the motions of the input test masses. We obtained the input-output relation without specifying the mirror displacements. The equations of motion of mirrors determine these mirror displacements. Since we discussed the quantum motion of mirrors, we used the Heisenberg equations of motion in quantum mechanics as the equations of motion for mirrors. Furthermore, we introduce the fundamental frequency $\omega_{p}$ of the mirrors’ pendulum, where $\omega_{p}$ is estimated as $\sim 1$Hz.
@@
In Sec. V, we briefly reviewed a quantum mechanical forced harmonic oscillator, which is essential to our arguments on the motions of mirrors in gravitational-wave detectors. Based on the understanding of this quantum forced harmonic oscillator, we explicitly derived the equations of motion for the mirrors in the Fabry-Përot interferometer in Sec. V.2. Then, we noted that evaluating the effects of radiation pressure is essential. Although the details are described in Appendix B, we review the outline of the derivation of the radiation pressure forces to the input test masses and to the end test masses in Fabry-Pérot gravitational-wave detectors in Sec. V.3. From the derived radiation pressure forces, we obtained the solutions to the Heisenberg equations in Sec. V.4.
@@
Together with the input-output relation derived in Sec. IV and the solutions to the Heisenberg equation derived in Sec. V.4, we obtain the final input-output relation of the Fabry-Pérot interferometer in Sec. VI. Then, we reach the input-output relation for the Fabry-Pérot interferometer (189). Although the parameters $\beta$, $\kappa$, and $h_{SQL}$ in the input-output relation are different from these parameters in Ref. [10], the resulting input-output relation is almost the same as that derived in Ref. [10]. The difference between $\kappa$ in this paper and the $\kappa$ in Ref. [10], which denote $\kappa^{(K)}$ in this paper are depicted in Fig. 3. The difference between $h_{SQL}^{(EQ)}$ ($m=m_{EM}=m_{ITM}$) and $h_{SQL}$ in Ref. [10] which denote $h_{SQL}^{(K)}$ within this paper are depicted in Fig. 4. The frequency dependence of these differences arises from the phase difference between the motion of EMs and that of ITMs. Due to these differences, we reach the difference of the $S_{(K)}^{(EQ)}$ and $S_{(K)}^{(K)}$ as depicted in Fig. 5, where $S_{(K)}^{(EQ)}$ is the equal mass version of Eq. (197) and $S_{(K)}^{(K)}$ is defined by Eq. (212).
@@
In addition to the above difference between the input-output relation (189), the difference also exists in the classical carrier part as the parameter $\mathfrak{R}$}}. This classical carrier field, modified by $\mathfrak{R}$}}, is used as the reference field to measure the gravitational-wave signal in the DC readout scheme. Therefore, the classical carrier field is essential in the DC readout scheme. As mentioned in Sec. VII, the expectation value of the signal is not affected by the additional parameter $\mathfrak{R}$}} in the classical carrier field due to the reality condition of the gravitational-wave signal $H(\Omega)=H^{*}(-\Omega)$. The contribution of $\mathfrak{R}$}} in the classical carrier field affects the signal-referred noise spectral density (234). For this reason, we discussed the difference from the Kimble noise spectral density in Ref. [10]. As a result of the additional classical carrier field $\mathfrak{R}$}}, we have to conclude that the shot noise contribution to the high-frequency range of the signal-referred noise spectral density does not decrease even if the incident laser power is increased. This contradicts the description in Ref. [10], which is the common knowledge in the community of gravitational-wave experiments.
@@
In this situation, we reconsidered the tuning points (161) and (177) in Sec. VIII. The origin of the additional modification $\mathfrak{R}$}} of the classical carrier in the input-output relation (189) is the leakage of the classical carrier from the radiation pressure forces, which are proportional to $2\pi\delta(\Omega)$ in the Fourier transformation of the radiation pressure forces. In the time domain, these terms, proportional to $2\pi\delta(\Omega)$, represent the constant forces in the radiation pressure forces. Here, we note that the equations of motion for mirrors are expressed as a forced harmonic oscillator. In simple classical mechanics, or even in quantum mechanics, a constant force in a harmonic oscillator yields the deviation of the equilibrium point. If there exists a constant force in the Heisenberg equation of motion for a harmonic oscillator, we have to take into account the change of the equilibrium points of the mirror displacements. For this reason, we reconsidered the Heisenberg equation of motion for the mirrors and the tuning points (161) and (177) in Sec. VIII.
@@
We also considered incomplete equilibrium tuning, in which deviations in the mirror displacements from their equilibrium points are partially eliminated through the introduction of the incomplete parameter $\epsilon$. We call this elimination the “incomplete equilibrium tuning.” If $\epsilon=0$, we can realize the complete equilibrium tuning discussed above. Instead of the complete equilibrium tuning (261) and (262), we apply the tuning condition (288) and (289). We also showed that the additional modification $\mathfrak{R}$}} of the classical carrier in the input-output relation is explicitly expressed by the displacement of the equilibrium points ${{\mathscr{D}}}_{EM}$ and ${{\mathscr{D}}}_{ITM}$. In this incomplete equilibrium tuning, the stationary noise spectral density is given by (292). To clarify the properties of the noise spectral density (292), we consider the typical two cases. The first case is $\theta\ll 1$ but $\theta\neq 0$ and the second case is $\theta=\pi/4$. These are depicted in Figs. 10 and 11.
In the first case, the noise spectral density is given by Eq. (294). If $\epsilon$ is sufficiently large, Fig. 10 indicates the dip in the low-frequency region. Although this dip violates the “standard quantum limit” which is the envelope of the curves with different powers $I_{0}$ in $\epsilon=0$ case, this is not surprising fact that there is no contradiction with the commutation relation $\hat{X}(t),\hat{P}(t)]=i\hbar$ as discussed in Sec. V. In the second case, the effect of $\mathfrak{R}$}} appears as the factor of the noise spectral density as Eq. (298). In both cases, if the incompleteness $\epsilon$ is sufficiently large, there are cases where the shot-noise contribution to the high-frequency region in the signal-referred noise spectral density does not decrease even if the incident power is increased. We also evaluate the incompleteness $\epsilon$ when we can realize the behavior near the complete equilibrium tuning in Eqs. (LABEL:eq:classical-terms-of-RPF-EM-renorm-retune-require-leading-order) and (LABEL:eq:classical-terms-of-RPF-ITM-renorm-retune-require-leading-order). This is the main result of this paper.
Here, we note that we cannot discuss the recent experimental results of the violation of the “standard quantum limit” [12, 13] through the ingredients of this paper, because we did not discuss the power recycling, the signal recycling, or the squeezed input techniques. We cannot say any relation between the dips depicted in Fig. 10 and works in Refs. [12, 13] due to the same reason. In other words, there is ample room for development through an accurate understanding of the ingredients in this paper. We leave these rooms for future work.
@@
Finally, we emphasize that the purpose of this paper is not to propose new techniques nor to develop new ideas for gravitational-wave detectors themselves. However, in this paper, we aim to explicitly show the derivation of the common knowledge in the community of gravitational-wave detections, thereby clarifying their concepts from a more theoretically accurate perspective. Due to the more precise theoretical arguments on the noise spectral density, we could discuss the incomplete equilibrium tuning of the Frabry-Pérot interferometer as an imperfection of the interferometric gravitational-wave detectors. We hope the ingredients of this paper and this kind of research will be helpful for the further development of detector science in gravitational-wave detectors.